Write the balanced equation for the combustion of isooctane (C8H18) to produce carbon dioxide and water. Use the smallest possible integers to balance the equation. Also, separate the + sign with 1 space.

2C8H18 + 25O2 ? 16CO2 + 18H2O

You are correct.
Your receipt no. is 161-2857 Help: Receipt
Previous Tries

Assuming gasoline is 90.0% isooctane, with a density of 0.692 g/mL, what is the theoretical yield (in grams) of CO2 produced by the combustion of 1.42 x 1010 gallons of gasoline (the estimated annual consumption of gasoline in the U.S.)? Remember, there are 3.785 liters in 1 gallon and assume that isooctane is the only carbon containing component of gasoline.

Scientific notation can be entered as follows: 1.23 x 1023 = 1.23E23

g

Compare the number of moles of C8H18 to the number of moles of CO2 from the balanced reaction.
Using this, the information given above, and the periodic table, set up an equation so the units cancel.


HINT: The balanced equation has 25O2
Answer #1
Concepts and reason

Combustion reaction is a reaction in which, a hydrocarbon reacts with oxygen to produce carbon dioxide and water.

Fundamentals

Mass, m of a substance is related to number of moles, n as follows:

m=n×Mm = n \times M

Here, M is molar mass of substance.

The balanced chemical equation of combustion of isooctane (C8H18)\left( {{{\rm{C}}_{\rm{8}}}{{\rm{H}}_{{\rm{18}}}}\;} \right) is

2C8H18+25O216CO2+18H2O2{{\rm{C}}_{\rm{8}}}{{\rm{H}}_{{\rm{18}}}}\; + {\rm{ }}25{{\rm{O}}_2}\;\longrightarrow{{}}{\rm{ }}16{\rm{C}}{{\rm{O}}_2}\; + {\rm{ }}18{{\rm{H}}_{\rm{2}}}{\rm{O}}

Volume of gasoline present (v)

v=(1.42×1010gallon)(3.785L1gallon)(1000mL1L)=5.37×1013mL\begin{array}{c}\\v = \left( {1.42 \times {{10}^{10}}{\rm{gallon}}} \right)\left( {\frac{{3.785{\rm{ L}}}}{{1{\rm{ gallon}}}}} \right)\left( {\frac{{1000{\rm{ mL}}}}{{1{\rm{ L}}}}} \right)\\\\ = 5.37 \times {10^{13}}{\rm{ mL}}\\\end{array}

Mass of gasoline (m)

m=dvm = dv

Where d is the density and v is the volume of gasoline

Substitute 0.69gmL1{\rm{0}}{\rm{.69 g m}}{{\rm{L}}^{ - 1}} for d and 5.37×1013mL5.37 \times {10^{13}}{\rm{ mL}} for v

m=(0.69gmL1)(5.37×1013mL)=3.7×1013g\begin{array}{c}\\m = \left( {{\rm{0}}{\rm{.69 g m}}{{\rm{L}}^{ - 1}}} \right)\left( {5.37 \times {{10}^{13}}{\rm{ mL}}} \right)\\\\ = 3.7 \times {10^{13}}{\rm{ g}}\\\end{array}

Now mass of isooctane (mi)\left( {{m_i}} \right) will be

mi=90100(m){m_i} = \frac{{90}}{{100}}\left( m \right)

Substitute 3.7×1013g3.7 \times {10^{13}}{\rm{ g}} for m

mi=90100(3.7×1013g)=3.34×1013g\begin{array}{c}\\{m_i} = \frac{{90}}{{100}}\left( {3.7 \times {{10}^{13}}{\rm{ g}}} \right)\\\\ = 3.34 \times {10^{13}}{\rm{ g}}\\\end{array}

Number of moles of isooctane

ni=miM{n_i} = \frac{{{m_i}}}{M}

Substitute 3.34×1013g3.34 \times {10^{13}}{\rm{ g}} for mi{m_i} and 114 g for M

ni=3.34×1013g114g=0.029×1013mol\begin{array}{c}\\{n_i} = \frac{{3.34 \times {{10}^{13}}{\rm{ g}}}}{{114{\rm{ g}}}}\\\\ = 0.029 \times {\rm{1}}{{\rm{0}}^{13}}{\rm{ mol}}\\\end{array}

According to equation, 2 moles isooctane gives 18 moles carbon dioxide

Hence 0.029 mol isooctane will give nCO2{n_{{\rm{C}}{{\rm{O}}_2}}} mol

nCO2=182(0.029×1013mol)=0.26×1013mol\begin{array}{c}\\{n_{{\rm{C}}{{\rm{O}}_2}}} = \frac{{18}}{2}\left( {0.029 \times {\rm{1}}{{\rm{0}}^{13}}{\rm{ mol}}} \right)\\\\ = 0.26 \times {\rm{1}}{{\rm{0}}^{13}}{\rm{ mol}}\\\end{array}

Hence theoretical yield of carbon dioxide in moles is 0.26×10130.26 \times {\rm{1}}{{\rm{0}}^{13}} mol

Theoretical yield of carbon dioxide in g (mCO2)\left( {{m_{{\rm{C}}{{\rm{O}}_2}}}} \right)

mCO2=nCO2×MCO2{m_{{\rm{C}}{{\rm{O}}_2}}} = {n_{{\rm{C}}{{\rm{O}}_2}}} \times {M_{{\rm{C}}{{\rm{O}}_2}}}

Substitute 0.26×10130.26 \times {\rm{1}}{{\rm{0}}^{13}} mol for nCO2{n_{{\rm{C}}{{\rm{O}}_2}}} and 44 g for MCO2{M_{{\rm{C}}{{\rm{O}}_2}}}

mCO2=0.26×1013mol×44g=11.6×1013g\begin{array}{c}\\{m_{{\rm{C}}{{\rm{O}}_2}}} = 0.26 \times {\rm{1}}{{\rm{0}}^{13}}{\rm{ mol}} \times 44{\rm{ g}}\\\\{\rm{ = 11}}{\rm{.6}} \times {\rm{1}}{{\rm{0}}^{13}}{\rm{ g}}\\\end{array}

According to equation, 2 moles isooctane gives 18 moles carbon dioxide

Hence ni{n_i} isooctane will give nCO2{n_{{\rm{C}}{{\rm{O}}_2}}} mol

nCO2=182(ni)=9ni\begin{array}{c}\\{n_{{\rm{C}}{{\rm{O}}_2}}} = \frac{{18}}{2}\left( {{n_i}} \right)\\\\ = 9{n_i}\\\end{array}

Ans:

Hence, the balanced equation is: 2C8H18+25O216CO2+18H2O2{{\rm{C}}_{\rm{8}}}{{\rm{H}}_{{\rm{18}}}}\; + {\rm{ }}25{{\rm{O}}_2}\;\longrightarrow{{}}{\rm{ }}16{\rm{C}}{{\rm{O}}_2}\; + {\rm{ }}18{{\rm{H}}_{\rm{2}}}{\rm{O}}

Theoretical yield of carbon dioxide is 11.6×1013g{\rm{11}}{\rm{.6}} \times {\rm{1}}{{\rm{0}}^{13}}{\rm{ g}}